Operation Summary for June 8, 2011 - June 10, 2011 |
|
Wednesday June 8 The day shift began with Tevatron (TEV) down due to a CAMAC link problem, with Operations monitoring the Recycler stash (196E10), with the Antiproton Source (Pbar) stacking (19.0E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off for target replacement, with the Meson FTBFT-978 experiment off due to high building temperatures, with the MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for beam. During the day shift, FESS personnel traced the loss of power that caused of the abort of store 8797 to the TEV sector A0 North main breaker power cables; they had switched to different cables by 11:51 AM. Unfortunately, the TEV quenched at A4 and the B0 low beta magnets due to a TCLK glitch. Operators prepared the TEV and D0 for access to perform vacuum work, inspect a portable sump pump, and cryo system maintenance. All accesses were completed by 1:41 PM. Operators put the TEV into recovery mode. At 2:51 PM, operators turned off MiniBooNE due to high service building temperatures. During the evening shift, Operations put the TEV into HEP shot setup and established store 8799 at 6:20 PM with an average initial luminosity (AIL) of 396.87E30. Operators reported that the MiniBooNE service building had cooled enough to run beam, which began at 7:14 PM. Thursday June 9 The midnight shift began with Operations monitoring store 8799 (~112.5E30) and the stash (175.4E10), with Pbar stacking (26.1E10), and with MiniBooNE taking beam. During the midnight shift, store 8799 aborted at 12:06 AM due to a lightning glitch. Operators prepared CDF for an access and when they completed their access turned the TEV back on. Operations put the TEV into shot setup and established store 8801 at 6 AM with a AIL of 311.73E30, which aborted at 7:20 AM due to another lightning strike. Operators switched off MI for an access to repair vacuum problems at 7:30 AM. During the day shift, Operations reported that the MI vacuum repairs lasted nearly all shift and they were successful. During this downtime, operators prepared the TEV A-E, F Sector, and Transfer Hall for access. FESS, TEV, and other personnel accessed the TEV to investigate vacuum problems, conduct other miscellaneous work, and to remove ground water and repair sumps. FESS personnel reported finding large amounts of water near E0; they believe the water got in because the E0 vehicle access ramp that runs down into the enclosure flooded. Operators removed the water and FESS personnel prepared for more rain by placing sand bags to keep water from reaching the tunnel through the access ramp. Additionally, Safety personnel accessed MI-12B to obtain sump samples and D0 personnel accessed their collision hall to conduct minor repairs. During the evening shift, operators began turning on MI and TEV. Operators found a problem with a MI quadrupole power supply (QF11) and called in EE Support personnel to make repairs. MiniBooNE started taking beam at 5:07 PM and Pbar resumed stacking at 6:06 PM. TEV personnel conducted Electron Cloud studies from 7 PM until 9 PM. Operators put the TEV through wet squeezes and orbit checks and then into shot setup at 10:34 PM. Operations established store 8804 at 11:39 PM with an AIL of 198.3E30. Friday June 10 The midnight shift began with Operations monitoring store 8804 (169.86E30) and the stash (44.0E10), with Pbar stacking (12.8E10), and with MiniBooNE taking beam. During the midnight shift, operators reported that Feeder #36 had a power glitch that caused a power outage in some of the experimental area Proton Line service buildings; FESS personnel restored the power. Operations established beam to T-978 at 4 AM. Operators reported that bad beam positions in MI restricted available intensities to T-978. The Plans for Friday and the Weekend The plans for today and this weekend are to stack, stash, and store. The Run Coordinator will allow MI personnel to conduct coalescing studies. NuMI is still on schedule to begin taking beam between 10 AM and noon tomorrow, 6/11/11. |
|
More Information Comments and Suggestions |
| last modified 06/10/2011 email Fermilab |
| Security, Privacy, Legal | |